II b. ASSUME the Differential Equation: dx dt -(wº)x. 1 SOLVE That is determine a function x = f(t) which satisfies the aboveplease help with b and c very neat and explain please! thanks
Answer #1

b) So we write differential as

\int \frac{dx}{x}=-\int \omega^2 dt

ln (x)=- \omega^2 t+C

So,

x(t)=Ce^{- \omega^2 t}

2)

So by diiffrential ths function we get

\frac{dx}{dt}=\frac{d}{dt}Ce^{- \omega^2 t}

so,

\frac{dx}{dt}=-\omega^2Ce^{- \omega^2 t}
so,

\frac{dx}{dt}=-\omega^2x

so this function satisfy our relation

c) So we write

\vec{A}.\vec{B}=(5\hat{i}+5\hat{j}+5\hat{k}).(\hat{j}-\hat{k})

\vec{A}.\vec{B}=5-5=0

2)

\vec{A}\times \vec{B}=(5\hat{i}+5\hat{j}+5\hat{k})\times(\hat{j}-\hat{k})

so,
\vec{A}\times \vec{B}=5\hat{k}-5\hat{i}+5\hat{j}-5\hat{i}

\vec{A}\times \vec{B}=5\hat{k}-10\hat{i}+5\hat{j}

3) So,

\vec{A}.(\vec{B}\times \vec{C})=\vec{C}.(\vec{A}\times \vec{B})

so,

=(-10\hat{i}).(5\hat{k}-10\hat{i}+5\hat{j})=100

and

D\vec{A}\times (\vec{C}\times \vec{B})=2(5\hat{i}+\5\hat{j}+5\hat{k})\times ((-10\hat{i})\times (\hat{j}-\hat{k}))

so,

D\vec{A}\times (\vec{C}\times \vec{B})=2(5\hat{i}+\5\hat{j}+5\hat{k})\times (-10\hat{k}-10\hat{j})

D\vec{A}\times (\vec{C}\times \vec{B})=2(50\hat{j}-50\hat{i}-50\hat{k}+50\hat{i})

D\vec{A}\times (\vec{C}\times \vec{B})=100\hat{j}-100\hat{k}